# CustomListener: Text-guided Responsive Interaction for User-friendly Listening Head Generation

### v1

@@ Abstract
Listening head generation aims to synthesize a non-verbal responsive listener head by modeling the correlation between the speaker and the listener in dynamic conversion. The applications of listener agent generation in virtual interaction have promoted many works achieving the diverse and fine-grained motion generation. However, they can only manipulate motions through simple emotional labels, but cannot freely control the listener’s motions. Since listener agents should have human-like attributes (e.g. identity, personality) which can be freely customized by users, this limits their realism. In this paper, we propose a user-friendly framework called CustomListener to realize the free-form text prior guided listener generation. To achieve speaker-listener coordination, we design a Static to Dynamic Portrait module (SDP), which interacts with speaker information to transform static text into dynamic portrait token with completion rhythm and amplitude information. To achieve coherence between segments, we design a Past Guided Generation Module (PGG) to maintain the consistency of customized listener attributes through the motion prior, and utilize a diffusion-based structure conditioned on the portrait token and the motion prior to realize the controllable generation. To train and evaluate our model, we have constructed two text-annotated listening head datasets based on ViCo and RealTalk, which provide text-video paired labels. Extensive experiments have verified the effectiveness of our model.
@@ 1 Introduction
Listening Head Generation (LHG) is to generate listener motions (face expressions, head movements, etc.) that respond synchronously to the speaker during face-to-face communication. Good communication is a two-way process, and the roles of speakers and listeners are equally important [4]. Whether the listener responds promptly will affect whether the speaker can actively express the follow-up content [17, 19].
Unlike speaker head generation (SHG) [7, 16, 31, 37], listeners can only convey information through non-verbal feedback (nodding, frowning, etc.), and their responses are affected by both sides of the interaction (speaker’s motion, tone, and listener’s own personality and emotion state). Currently, many applications, such as digital avatar generation [8, 20] and human-computer interaction [42], use the listener agent to simulate more realistic conversations, which has promoted many studies in LHG.
@@
Although above methods can combine speaker-listener information to generate diverse and fine-grained listener motions, they have restrictions in the explicit control of the listener agent’s response. Specifically, L2L [25] focuses on the motion diversity but cannot control the motion under specific conditions. RLHG [43], PCH [15], MFR-Net [21], and ELP [33] achieve controllable listener motion conditioned on the input attitude. However, the attitude is limited to a few fixed labels (eg. positive, neutral, negative), posing limitations on the generation of realistic listeners: First, simple labels are not enough to give accurate responses. For example, for negative attitudes, the two emotions of doubt and anger correspond to different facial expressions. More importantly, a realistic listener agent should have its own identity (parent, friend, teacher, etc.), personality (lively, calm, etc.) and behavioral habits (e.g. frowning when thinking), which users can customize and set in advance. However, simple labels cannot achieve the control under these conditions.
Considering above limitations of simple labels, in this paper, we propose a user-friendly framework called CustomListener to enable the text-guided generation, as shown in Figure 1. Users can pre-customize detailed attributes of the listener agent. Then using a large language model, we combine speech content and user-customized attributes to obtain the text prior of the listener’s basic portrait. In CustomListener, we seamlessly incorporate speaker information while being guided by the text prior, so as to generate realistic listener responses that are controllable and interactive.
@@
To solve above challenges, in our CustomListener, we design the Static to Dynamic Portrait module (SDP) and the Past-guided Motion Generation module (PGG) to achieve the coordination and the coherence respectively. In SDP, we transform the static portrait provided by the text prior to be dynamic one, which interacts with audio to obtain the time-dependent information about the completion time and rhythm of the gradual motion changes, and refine the listener’s motions to make the speaker-listener fluctuations relevant. In PGG, we generate a motion prior based on the similarity of dynamic portrait tokens between adjacent segments, which provides reference information for features that need to be maintained in the current compared to the past. Then, the motion prior and the dynamic portrait token are input as conditions into the diffusion-based structure [14] with fixed noise at the segment connection, finally generating listener coefficients that enable coordination and coherence in interactions.
@@
  • For long-term generation, we design a PGG module to meet the listener motion coherence between different text-prior-conditioned segments, so that the consistency of the listener’s customized behavioral habits between segments can be maintained and the motion switching is smooth.
@@ 2 Related work
Figure 2: Overall framework of CustomListener. Given the text prior providing the listener’s static portrait style, SDP-Module transforms the static portrait into a dynamic one, which includes the time-dependent information and the fluctuation of motions. Then in PGG-Module, two adjacent portrait tokens are input to Past-guided Module to generate the motion prior, which is helpful to maintain the motion coherence. The portrait token and the motion prior are utilized as conditions of the diffusion-based structure to realize the controllable generation.
@@
Some early works [34, 1, 5] use rule-based methods to manually incorporate interactive motions, which have limitations in motion diversity and generalization on the wild data. Later, data-driven methods [11, 26] generate 2D motions based on facial keypoints or movement frequencies, but 2D information cannot fully represent the details of facial movements.
Recently, motion generation based on 3D face coefficients, coupled with the post-processing renderer, has gradually become mainstream due to its stronger representation ability of facial movements. RLHG [43] proposes a representative ViCo dataset and constructs a baseline coefficient decoding method using a sequence-sequence structure. After that, PCH [15] assembles an enhanced renderer to make the generated results visually better and more stable by inpainting the image boundary and background. Learning2Listen (L2L) [25] proposes to use a sequence-encoding VQ-VAE [35] to learn the discrete codebook of listener motion. ELP [33] further utilizes a composition of multiple discrete motion-codewords to represent facial motion in a more fine-grained manner.
MFR-Net [21] achieves the preservation of listener’s identity and the diversity of generation based on the self-designed feature aggregation module.
@@
The motion generation task derives applications under a variety of conditions. For example, body motion conditioned on text descriptions [18, 27, 40, 28, 13], motion trajectory conditioned on scene image [3], and dance motion conditioned on music [38]. Recently, large language models promote the development of free-form text conditioned generation based on auto-encoder [28], hierarchical model [13], and diffusion models [18]. We also generate motions conditioned on free-form text. However, for our responsive interaction, in addition to focusing on the matching of motion and text like above methods, we also need to make the motion and the speaker’s action, speech, tone and rhythm harmonious to express the listener’s empathy. In our paper, we adopt the SDP module to achieve this harmony. For long-term generation, to ensure the coherent fluency between clips under different text commands, some methods have adopted transition sampling [39], specifying noise [41], etc.. However, the coherence for the custom listener, which has some motion habits, is more complex. It is not just a simple motion switch, but is related to the past habits and styles, and our PGG module solve this challenge.
@@ 3.1 Overview
In our CustomListener framework, under the guidance of the fine-grained text prior, we can synthesize natural and controllable listener motions, which can seamlessly integrate the speaker’s information (motion, speech content, audio) to give interactive non-verbal responses. Let $L_{m}\in\mathbb{R}^{T\times 70}$ denote the listener motions to be generated, and $T$ is the number of frames, 70 denotes 64-dim expression and 6-dim pose. $R$ denotes the text prior. Given speaker’s motions $S_{m}\in\mathbb{R}^{T\times 70}$, audio $S_{a}$ and speech content $S_{t}$, our framework can be formulated as:
$L_{m}=Custom(S_{m},S_{a},S_{t};R)$ (1)
The overview of our proposed framework is shown in Figure 2. Before generation, for each video segment, we use GPT to generate a text prior based on user-customized listener attributes and the speaker’s speech content, which provides the listener’s static portrait reference. Then, we extract speaker’s audio features from the Audio Encoder and speaker’s motion coefficients from a pre-trained face reconstruction model [9]. Following this, we generate realistic listener motions through a two-stage process: (1) Static to Dynamic Portrait generation (SDP-Module): it can transform static portrait tokens into the dynamic through audio-text responsive interaction and speaker-motion based refinement. The dynamic portrait token includes the time-dependent information about the motion completion and rhythm, and makes the speaker-listener motion fluctuations correlated; (2) Past-guided Listener Motion Generation (PGG-Moudle): to produce coherent motions and maintaining the listener’s previous behavioral habits in long-term generation, we use a past guided module to produce the motion prior, combined with the diffusion denoising process with a fixed initial noise to generate the listener coefficients. Finally, listener videos are obtained through the post-rendering process.
@@ 3.2 Static to Dynamic Portrait generation
In this section, we first introduce the generation of the static portrait-token, and subsequently, we present SDP-Module, which can convert static tokens into dynamic one by two stages: (1) Audio-text Responsive Interaction: static portrait tokens are converted into time-dependent portrait tokens, which reflects the completion time and rhythm of motions according to the semantics and intonation in the audio, ensuring the gradual change of the motion; (2) Speaker Motion-based Refinement: it refines the amplitude of the listener’s motions according to the fluctuations of speaker’s motions, producing the final dynamic portrait token combined with the time-dependent portrait token.
@@
To establish a better connection between the user-customized listener attributes (identity, personality, relationship, attitude, etc.) and the listener’s motions, we leverage the powerful GPT for semantic understanding of both user-customized contents and speaker’s speech contents, and generate the text prior that accurately describes the listener’s expressions and poses. Following this, we use RoBERTa [22], a robust and efficient model with high natural language understanding capability, to encode the text prior into text embeddings. To obtain portrait-related tokens, we further employ a mapping net to convert text embeddings into portrait tokens. Since the text prior describes the listener’s static portrait including expressions and head movements in the entire video clip, we consider the above portrait tokens generated from the text prior to be static portrait tokens.
@@
Since static portrait tokens from text prior provide a comprehensive representation of listener head in a video segment, it lacks the time-dependent information when the motion is completed. Thus, the model would be limited to generating constant motion and cannot effectively convey gradual variation in movement without a specific design. To address this critical issue, we make use of the semantic knowledge in speaker’s utterances, which serves as a guidance for motion shifts. Specifically, we choose the speaker audio to provide semantics instead of speech contents. The underlying intuition is that the former contains not only semantic information but also the temporal information. This entire process can be splited into two parts, part A and part B.
Figure 3: Illustration of Audio-text Responsive Interaction. We first generate a weight matrix by responsive interaction between audio features and static portrait-token (SP Token), and then generate time-dependent portrait tokens (TP Token) guided by this weight matrix.
@@
Part B After obtaining a weight matrix $M(A,H)$, we can generate time-dependent portrait tokens based on static portrait tokens with the guidance of $M(A,H)$. Specifically, as illustrated in Figure 3, the time-dependent portrait tokens (TP Token) at time stamp $k$ is generated with the highest weighting assigned to the static portrait tokens (SP Token) at $l=j$, which corresponds to the phrase “raised lip corner”. This part can be formulated as Equation 3.
@@
where $Sm$ denotes Softmax, $A\in\mathbb{R}^{T\times C}$ is the audio feature, $H\in\mathbb{R}^{L\times C}$ is the static portrait token, $M\in\mathbb{R}^{T\times L}$ is the weight map, $\alpha$ is a scaling factor, $E_{t}\in\mathbb{R}^{1\times C}$ is the time-dependent portrait token at time stamp $t$. $L$ is the number of word segments in the text, and $C$ is the feature dimension.
@@
In human-to-human conversations, we believe that the amplitude of the speaker’s motion fluctuation also has an impact on the changes in the listener’s motion response. Therefore, we use the cumulative value of the speaker’s motion changes in the past window period to refine the listener’s motion amplitude. This process can be well-designed formulated as:
@@
where $\oplus$ means concatenation, $E_{t},\overline{E_{t}},\widetilde{E_{t}}$ is time-dependent portrait token, speaker motion-weighted portrait token and dynamic portrait token at frame $t$, respectively. $S_{i}$ denote speaker motion at frame $i$, $w$ is the time-window length. For efficiency, we set $w$ to 5, which means we only take the past five-frame speaker motions into account. For each time stamp $t$, the listener’s portrait token $E_{t}$ is affected by the cumulative gradients of several previous speaker motions. This process can be illustrated in the lower part of the SDP-Module in Figure 2. Notably, when generating portrait token at frame $t$, we only focus on the speaker motions up to frame $t-1$. Due to the listener and the speaker communicating concurrently, the listener’s state at frame $i$ should not influenced by the speaker’s motion at the same frame $i$.
@@ 3.3 Past-guided Listenr Motion Generation
The PGG-Module is to ensure that motions between segments are coherent, and the customized listener attributes reflected in the previous segment can be maintained in the current segment. It contains Past-guided Module and diffusion-based motion generation module, which can generate motion prior conditioned on past motions and then synthesize natural listener motions. Following this, we employ a pretrained render model [30] to obtain final listener videos.
@@
To produce long video with arbitrary length, we opt for a segment-by-segment generation approach during inference. However, we have observed that there are unsmooth pose transitions between segments. This issue can be attributed to the neglect of correlations between adjacent segments. To address this, we take past motions into consideration. Since our final goal is to synthesize a customized listener, in addition to ensuring smooth transition, we also need to take the listener’s past behavioral habits into account (e.g. frowning when thinking). Hence, we sought assistance from dynamic portrait tokens of two adjacent segments to guide the generation of motion prior.
Specifically, the weight we assign to past motions at different times is determined by the similarity between portrait tokens in the current segment and those in the last segment. Formally, we calculate the motion weight and obtain the motion tokens by:
@@
$G_{t:t+L}=Sim_{t:t+L}\times m_{t-L:t},$ (7)
where $\times$ denotes matrix multiplication, $L$ is the segment length, $t$ is the time stamp, $G_{t:t+L}$ is the generated motion prior, $Sim_{t:t+L}$ denotes the similarity between portrait tokens in current segment $E_{t:t+L}$ and those in last segment $E_{t-L:t}$, which is the weight we assign to past motions $m_{t-L:t}$. The higher the similarity, the greater the weight we assign. This approach can not only encourage continuous transitions between two adjacent segments, but also ensure the maintenance of customized listener’s behavioral habits between two adjacent sequences.
@@
To achieve diverse and natural head motion generation, our generation module is built on denoising diffusion probabilistic models (DDPM) [14]. And thus we can fix the initial noise of current segment to be the end of the previous segment to ensure smooth motion switching.
@@
where $c$ is a conditioning variable. We perform denoising process with a transformer decoder [36] conditioned on the motion prior and dynamic portrait token.

### v2

@@ Abstract
Listening head generation aims to synthesize a non-verbal responsive listener head by modeling the correlation between the speaker and the listener in dynamic conversion. The applications of listener agent generation in virtual interaction have promoted many works achieving diverse and fine-grained motion generation. However, they can only manipulate motions through simple emotional labels, but cannot freely control the listener’s motions. Since listener agents should have human-like attributes (e.g. identity, personality) which can be freely customized by users, this limits their realism. In this paper, we propose a user-friendly framework called CustomListener to realize the free-form text prior guided listener generation. To achieve speaker-listener coordination, we design a Static to Dynamic Portrait module (SDP), which interacts with speaker information to transform static text into dynamic portrait token with completion rhythm and amplitude information. To achieve coherence between segments, we design a Past Guided Generation module (PGG) to maintain the consistency of customized listener attributes through the motion prior, and utilize a diffusion-based structure conditioned on the portrait token and the motion prior to realize the controllable generation. To train and evaluate our model, we have constructed two text-annotated listening head datasets based on ViCo and RealTalk, which provide text-video paired labels. Extensive experiments have verified the effectiveness of our model.
@@ 1 Introduction
Listening Head Generation (LHG) is to generate listener motions (face expressions, head movements, etc.) that respond synchronously to the speaker during face-to-face communication. Good communication is a two-way process, and the roles of speakers and listeners are equally important [4]. Whether the listener responds promptly will affect whether the speaker can actively express the follow-up content [17, 19]. Unlike speaker head generation (SHG) [7, 16, 31, 39], listeners can only convey information through non-verbal feedback (nodding, frowning, etc.), and their responses are affected by both sides of the interaction (speaker’s motion, tone, and listener’s own personality and emotion state). Currently, many applications, such as digital avatar generation [8, 20] and human-computer interaction [45], use the listener agent to simulate more realistic conversations, which has promoted many studies in LHG.
@@
Although the above methods can combine speaker-listener information to generate diverse and fine-grained listener motions, they have restrictions in the explicit control of the listener agent’s response. Specifically, L2L [25] focuses on the motion diversity but cannot control the motion under specific conditions. RLHG [46], PCH [15], MFR-Net [21], and ELP [35] achieves controllable listener motion conditioned on the input attitude. However, the attitude is limited to a few fixed labels (eg. positive, neutral, negative), posing limitations on the generation of realistic listeners: First, simple labels are not enough to give accurate responses. For example, for negative attitudes, the two emotions of doubt and anger correspond to different facial expressions. More importantly, a realistic listener agent should have its own identity (parent, friend, teacher, etc.), personality (lively, calm, etc.) and behavioral habits (e.g. frowning when thinking), which users can customize and set in advance. However, simple labels cannot achieve the control under these conditions.
Considering the above limitations of simple labels, in this paper, we propose a user-friendly framework called CustomListener to enable the text-guided generation, as shown in Figure 1. Users can pre-customize detailed attributes of the listener agent. Then using a large language model, we combine speech content and user-customized attributes to obtain the text prior of the listener’s basic portrait. In CustomListener, we seamlessly incorporate speaker information while being guided by the text prior, so as to generate realistic listener responses that are controllable and interactive.
@@
To solve the above challenges, in our CustomListener, we design the Static to Dynamic Portrait module (SDP) and the Past Guided Generation module (PGG) to achieve the coordination and the coherence respectively. In SDP, we transform the static portrait provided by the text prior to be dynamic one, which interacts with audio to obtain the time-dependent information about the completion time and rhythm of the gradual motion changes, and refines the listener’s motions to make the speaker-listener fluctuations relevant. In PGG, we generate a motion prior based on the similarity of dynamic portrait tokens between adjacent segments, which provides reference information for features that need to be maintained in the current compared to the past. Then, the motion prior and the dynamic portrait token are input as conditions into the diffusion-based structure [14] with fixed noise at the segment connection, finally generating listener coefficients that enable coordination and coherence in interactions.
@@
  • For long-term generation, we design a PGG module to meet listener motion coherence between different text-prior-conditioned segments, so that the consistency of the listener’s customized behavioral habits between segments can be maintained and the motion switching is smooth.
@@ 2 Related work
Figure 2: Overall framework of CustomListener. Given the text prior providing the listener’s static portrait style, SDP-Module transforms the static portrait into a dynamic one. Then in PGG-Module, the dynamic portrait token are combined with motion prior generated from Past-guided Module and are utilized as conditions of the diffusion-based structure to realize the controllable generation. The ’C’ in the figure denotes concatenation and the small pink squares denotes diffusion time-step token.
@@
Some early works [36, 1, 5] use rule-based methods to manually incorporate interactive motions. Later, data-driven methods [11, 26] generate 2D motions based on facial keypoints or movement frequencies, but 2D information cannot fully represent the details of facial movements.
Recently, motion generation based on 3D face coefficients, coupled with the post-processing renderer, has gradually become mainstream due to its stronger representation ability of facial movements. RLHG [46] proposes a representative ViCo dataset and constructs a baseline coefficient decoding method using a sequence-sequence structure. After that, PCH [15] assembles an enhanced renderer to make the generated results visually better and more stable by inpainting the image boundary and background. Learning2Listen (L2L) [25] proposes to use a sequence-encoding VQ-VAE [37] to learn the discrete codebook of listener motion. ELP [35] further utilizes a composition of multiple discrete motion-codewords to represent facial motion in a more fine-grained manner. MFR-Net [21] achieves the preservation of listener’s identity and the diversity of generation.
@@
The motion generation task derives applications under a variety of conditions. For example, body motion conditioned on text descriptions [18, 27, 43, 28, 13], motion trajectory conditioned on scene image [3], and dance motion conditioned on music [41]. Recently, large language models promote the development of free-form text conditioned generation based on auto-encoder [28], hierarchical model [13], and diffusion models [18]. We also generate motions conditioned on free-form text. However, for our responsive interaction, in addition to focusing on the matching of motion and text like the above methods, we also need to make the motion and the speaker’s action, speech, tone and rhythm harmonious to express the listener’s empathy. In our paper, we adopt the SDP module to achieve this harmony. For long-term generation, to ensure the coherent fluency between clips under different text commands, some methods have adopted transition sampling [42], specifying noise [44], etc. However, the coherence for the custom listener, which has some motion habits, is more complex. It is not just a simple motion switch, but is related to past habits and styles, and our PGG module solves this challenge.
@@ 3.1 Overview
In our CustomListener framework, under the guidance of the fine-grained text prior, we can synthesize natural and controllable listener motions, which can seamlessly integrate the speaker’s information (motion, speech content, audio) to give interactive non-verbal responses. Let $L_{m}\in\mathbb{R}^{T\times 70}$ denote the listener motions to be generated, and $T$ is the number of frames, 70 denotes 64-dim expression and 6-dim pose. Given speaker’s motions $S_{m}\in\mathbb{R}^{T\times 70}$, audio $S_{a}$, speech content $S_{c}$ and text prior $R$, our framework can be formulated as:
$L_{m}=Custom(S_{m},S_{a},S_{c};R)\vspace{-0.2cm}$ (1)
The overview of our proposed framework is shown in Figure 2. Before generation, for each video segment, we use GPT to generate a text prior based on user-customized listener attributes and the speaker’s speech content, which provides the listener’s static portrait reference. Then, we extract the speaker’s audio features from the Audio Encoder and speaker’s motion coefficients from a pre-trained face reconstruction model [9]. Following this, we generate realistic listener motions through a two-stage process: (1) Static to Dynamic Portrait generation (SDP-Module): it can transform static portrait tokens into the dynamic through audio-text responsive interaction and speaker-motion based refinement. The dynamic portrait token includes the time-dependent information about the motion completion and rhythm, and makes the speaker-listener motion fluctuations correlated; (2) Past-guided Listener Motion Generation (PGG-Module): to produce coherent motions and maintain the listener’s previous behavioral habits in long-term generation, we use a past guided module to produce the motion prior, combined with the diffusion denoising process with a fixed initial noise to generate the listener coefficients. Finally, listener videos are obtained through the post-rendering process.
@@ 3.2 Static to Dynamic Portrait generation
In this section, we first introduce the generation of the static portrait-token, and then present SDP-Module, which can convert static tokens into dynamic one by two stages: (1) Audio-text Responsive Interaction: static portrait tokens are converted into time-dependent portrait tokens, which reflects the completion time and rhythm of motions according to the semantics and intonation in the audio, ensuring gradual motion changes; (2) Speaker Motion-based Refinement: it refines the listener’s motions according to the fluctuations of speaker’s motions, producing the final dynamic portrait token combined with the time-dependent portrait token.
@@
To establish a better connection between the user-customized listener attributes (identity, personality, relationship, attitude, etc.) and the listener’s motions, we leverage the powerful GPT for semantic understanding of both user-customized contents and speaker’s speech contents, and generate the text prior that accurately describes the listener’s expressions and poses. Following this, we use RoBERTa [22], a robust and efficient model with high natural language understanding capability, to encode the text prior into text embeddings. To obtain portrait-related tokens, we further employ a mapping net to convert text embeddings into portrait tokens. We consider these portrait tokens to be static portrait tokens since the text prior describes the listener’s static portrait (eg. expressions and poses) in the entire video clip.
@@
Since static portrait tokens from text prior provide a comprehensive representation of listener head in a video segment, it lacks the time-dependent information. Thus, the model would be limited to generating constant motion and cannot effectively convey gradual variation in movement without a specific design. To address this critical issue, we make use of the semantic knowledge in speaker’s utterances, which serves as a guidance for motion shifts. Specifically, we choose the speaker audio to provide semantics instead of speech contents. The underlying intuition is the former contains not only semantic information but also the temporal information. This entire process can be split into two parts, part A and part B.
Figure 3: Illustration of Audio-text Responsive Interaction. We first generate a weight matrix by responsive interaction between audio features and static portrait-token (SP Token), and then generate time-dependent portrait token (TP Token) guided by weight matrix.
@@
Part B After obtaining a weight matrix $M(A,H)$, we can generate time-dependent portrait tokens based on static portrait tokens with the guidance of $M(A,H)$. Specifically, as illustrated in Figure 3, the time-dependent portrait tokens (TP Token) at the time stamp $k$ is generated with the highest weighting assigned to the static portrait tokens (SP Token) at $l=j$, which corresponds to the phrase “raised lip corner”. This part can be formulated as Equation 3.
@@
where $Sm$ denotes Softmax, $A\in\mathbb{R}^{T\times C}$ is the audio feature, $H\in\mathbb{R}^{L\times C}$ is the static portrait token, $M\in\mathbb{R}^{T\times L}$ is the weight map, $\alpha$ is a scaling factor, $E_{t}\in\mathbb{R}^{1\times C}$ is the time-dependent portrait token at time stamp $t$. $L$ is the length of text embeddings, and $C$ is the feature dimension.
@@
In human-to-human conversations, we believe that the amplitude of the speaker’s motion fluctuation also has an impact on the changes in the listener’s motion response[32, 21]. Therefore, we utilize the past speaker’s motion changes to refine the listener’s motion amplitude. This process can be well-designed formulated as:
@@
where $\oplus$ means concatenation, $E_{t},\overline{E_{t}},\widetilde{E_{t}}$ is time-dependent portrait token, speaker motion-weighted portrait token and dynamic portrait token at frame $t$, respectively. $S_{i}$ denotes speaker motion at frame $i$, $w$ is the time-window length. For efficiency, we set $w$ to 5, which means we only take the past five-frame speaker motions into account. For each time stamp $t$, the listener’s portrait token $E_{t}$ is affected by the cumulative gradients of several previous speaker motions. This process can be illustrated in the lower part of the SDP-Module in Figure 2. Notably, when generating portrait token at frame $t$, we only focus on the speaker motions up to frame $t-1$. Due to the listener and the speaker communicating concurrently, the listener’s state at frame $t$ should not influenced by the speaker’s motion at the same frame $t$.
@@ 3.3 Past-guided Listenr Motion Generation
The PGG-Module is to ensure that motions between segments are coherent, and the customized listener attributes in the previous segment can be maintained in the current segment. It contains Past-guided Module and diffusion-based motion generation module, which can generate motion prior based on past motions and then synthesize listener motions. Following this, a renderer is used to obtain listener videos.
@@
To produce long video with arbitrary length, we opt for a segment-by-segment generation approach during inference. However, we have observed that there are unsmooth pose transitions between segments. This issue can be attributed to the neglect of correlations between adjacent segments. To address this, we take past motions into consideration.
Since our final goal is to synthesize a customized listener, in addition to ensuring smooth transition, we also need to take the listener’s past behavioral habits into account (e.g. frowning when thinking). Hence, we sought assistance from dynamic portrait tokens of two adjacent segments to guide the generation of motion prior. Specifically, the weight we assign to past motions at different times is determined by the similarity between portrait tokens in the current segment and those in the last segment. Formally, we calculate the motion weight and obtain the motion prior $G_{t:t+L}$ by:
@@
$G_{t:t+L}=Sim_{t:t+L}\times P_{t-L:t},$ (7)
where $\times$ denotes matrix multiplication, $L$ is the segment length, $t$ is the time stamp, $Sim_{t:t+L}$ denotes the similarity between portrait tokens in the current segment $E_{t:t+L}$ and those in the last segment $E_{t-L:t}$, which is the weight we assign to past motions $P_{t-L:t}$. The higher the similarity, the greater the weight we assign. This approach can not only encourage continuous transitions between two adjacent segments, but also ensure the maintenance of customized listener’s behavioral habits between two adjacent sequences.
@@
To achieve diverse and natural head motion generation, our generation module is built on denoising diffusion probabilistic models (DDPM) [14].
@@
where $c$ is a conditioning variable. We perform denoising process with a transformer decoder [38] conditioned on the motion prior and dynamic portrait tokens.